Write an equation of a line that is perpendicular to y= -3/4x + 2 and passes through (-4,1)
Answer by checkley71(8403)

THIS LINE HAS A SLOPE OF -3/4.
THEREFORE A PERPENDICULAR LINE HAS A SLOPE OF 4/3.
NOW REOPLACE THE X & Y VALUES IN THE LINE EQUATION [Y=mX+b] WITH (-4,1) & SOLVE FOR THE Y INTERCEPT (b).
1=4/3*-4+b
1=-16/3+b
b=1+16/3
b=(3+16)/3
b=19/3 FOR THE Y INTERCEPT.
Y=4X/3+19/3 ANSWER.